# The Remarkable Robustness of LLMs: Stages of Inference?

### v1

@@ Abstract
We demonstrate and investigate the remarkable robustness of Large Language Models by deleting and swapping adjacent layers. We find that deleting and swapping interventions retain 72-95% of the original model’s prediction accuracy without fine-tuning, whereas models with more layers exhibit more robustness. Based on the results of the layer-wise intervention and further experiments, we hypothesize the existence of four universal stages of inference across eight different models: detokenization, feature engineering, prediction ensembling, and residual sharpening. The first stage integrates local information, lifting raw token representations into higher-level contextual representations. Next is the iterative refinement of task and entity-specific features. Then, the second half of the model begins with a phase transition, where hidden representations align more with the vocabulary space due to specialized model components. Finally, the last layer sharpens the following token distribution by eliminating obsolete features that add noise to the prediction.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities, often attributed to their increased scale [67]. However, the benefits of scaling are accompanied by heightened risks and vulnerabilities [7, 36, 4], necessitating extensive research into the underlying mechanisms of these capabilities. Inspired by previous studies on model robustness [28, 45, 58, 44, 5, 65], this work investigates the sensitivity of LLMs to the deletion and swapping of entire layers during inference. Our findings suggest four universal stages of inference: detokenization, feature engineering, prediction ensembling, and residual sharpening.
Recent work in mechanistic interpretability has explored the iterative inference hypothesis [3, 58], which suggests that each layer incrementally updates the hidden state of a token in a direction of decreasing loss by gradually shaping the next token distribution [24]. Self-repair [58] and redundancy [45, 28] in networks further support this hypothesis of iterative inference. However, recent work also indicates a degree of specialization in networks, with attention heads and neurons playing specific roles [32, 43, 26], which compose into more sophisticated circuits [52, 21].
In this work, we begin by exploring the robustness of language models by performing a series of interventions that delete individual layers or swap adjacent layers (Figure 2). Using these results, we then attempt to understand the roles of different depths in the network. Our experiments suggest four phases in a model, which we investigate further.
Specifically, we hypothesize an initial (1) detokenization [15] stage, where the model integrates local context to convert raw token representations into coherent entities, as suggested by the sensitivity to deletion and swapping. In the (2) feature engineering stage, the model iteratively builds feature representations based on token context, leading to minimal progress in token prediction but significant increases in probing accuracy and patching importance. A phase transition from attention-heavy computation to MLP-heavy computation delineates the following stage. During the (3) prediction ensembling stage, the model emphasizes relevant predictions while suppressing others, potentially marked by high MLP computation and the emergence of prediction neurons. Finally, in the (4) residual sharpening stage, tokens transition from semantic representations to specific next-token predictions, again showing sensitivity to deletion and swapping. This framework represents a tentative step toward understanding the complex nature of token processing in advanced language models (see Table 1).
Figure 1 delineates our four characteristic phases, which we describe further in Table 1. We integrate our work with previous depth-related findings within interpretability [21].
@@
| 1 | Detokenization | Integrate local context to transform raw token representations into coherent entities | Catastrophic sensitivity to deletion and swapping |
@@
| 2 | Feature Engineering | Iteratively build feature representation depending on token context | Little progress made towards next token prediction, but significant increase in probing accuracy and patching importance. Attention Heavy Computation |
@@
| 3 | Prediction Ensembling | Convert previously constructed semantic features into plausible next token predictions using an ensemble of model components. | Increased MLP importance; prediction neurons appear; phase transition in progress towards final prediction |
@@
| 4 | Residual Sharpening | Sharpen the next token distribution by eliminating obsolete features that add noise to the prediction | More suppression neurons than prediction neurons |
2 Related Work
Universal Mechanisms
A key activity of mechanistic interpretability is circuit analysis, where research uncovers relevant model components for a given computation. In computer vision, circuits discover how features are constructed across many layers [51]. Follow-up work found that feature building was carried out by specific mechanisms that appeared across models, such as frequency detectors [60] and curve-circuits [8]. Language models seem to be following a similar line of inquiry, first uncovering universal model components, such as induction heads [52], successor heads [26], and copy suppression [43] in attention mechanisms. The discovery of knowledge neurons [10] paved the way for the identification of various specialized neurons [32, 66]. These specialized components can be connected to critical roles in universal processes in language models, such as circuit reuse [47], variable finding mechanisms [19], self-repair [58, 44] (which also studies layer-wise ablations), function vectors [63, 35], and long context retrieval [64].
Depth-Dependent Studies
Circuit analysis naturally motivated various depth-dependent studies, such as the logit lens [50], a technique that reveals the model’s prediction distribution at each layer. This line of reasoning revived the iterative inference hypothesis, the idea that each layer updates the hidden state in a direction of decreasing loss, in the context of ResNets [27, 37]. Due to residual connections, researchers were able to demonstrate that models exhibited “ensembling" [65] through layer ablations and permutations, referred to as “lesion" studies. This was later applied to modern transformers [13, 5]. Recent work [3] expanded on the logit lens and provided further evidence for iterative inference. This hypothesis was also supported by analyzing transformers in embedding spaces [12] and model self-repair [58].
Many works have localized specific kinds of computations to particular regions of large language models. For instance, model editing research [46] showed that knowledge is stored in mid-layer MLP neurons. Follow-up studies [33, 25] suggested that these facts can be stored across layers, with different components encoding task-specific instructions and recall. There is significant literature citing changes halfway through a language model, such as linear probe quality improving the fastest in the first half of a model [30] or fine-tuning predominantly updating weights in the middle of a model [53]. Few works observed that activation sparsity changes in the middle of a model, transitioning from sparse to dense [40, 66].
Other works have extended these insights into phases, such as identifying emergent specific phases of truth processing in language models [41] and distilling translation in multilingual transformers into three distinct phases: input space, concept space, and output space [68]. Other works hypothesized stages of inference in large language models [15]. This was further studied in the context of iterative inference [3], which demonstrated the importance of the first layers through layer ablations. When permuting the layers of GPT-2 style transformers, studies found the best performance in variants that have a higher proportion of self-attention layers at the beginning of models and more feedforward layers at the end [54]. These studies thematically suggest a preferential order to computation, which we investigate further.
Robustness: Quantization, Pruning, Redundancy
Numerous studies have inadvertently revealed depth-related findings while exploring model pruning or quantization. Prior work on zero-ablations of transformers has predominantly concentrated on BERT [14] style transformers [71, 17, 18, 59, 69]. Despite significant differences in these models [16], many underlying principles may be applicable. For example, pruning the final layers of a model retains most of the model’s performance [59]. Follow-up studies suggest 85% neuron redundancy in BERT transformers [11]. This is analogous to recent findings in modern networks, which demonstrate that approximately 70% of attention heads and 20% of feed-forward networks can be removed with minimal impact on task performance [1], suggesting redundancy [28, 45]. Model quantization work found improved benchmark performance by only keeping low-rank components of MLPs in the second half of the models [61] while uncovering token frequency dependencies in MLP weights. While we only touch upon the relationship between robustness and token frequency, [42] proposes that pretraining data contributes to this robustness.
@@ 3 Experimental Protocol
To investigate the stages of inference in language models, we examine the Pythia [6], GPT-2 [56], and Microsoft Phi [29, 39] model families, which range from 124M to 6.9B parameters (see Table 2). All families use decoder-only transformers but exhibit several architectural differences that enable us to test the generality of our findings. Pythia models utilize parallel attention and MLPs, executing these components concurrently during inference. In contrast, GPT-2 and Phi models apply these components sequentially, with attention followed by the MLP (see Figure 2). Additionally, GPT-2 models were trained with dropout. We preprocess weights identically across all models, folding in the layer norm, centering the unembedding weights, and centering the writing weight. Despite the architectural differences, these models exhibit consistent inference patterns that validate our hypothesis.
@@
We evaluate all three model families on a corpus of one million tokens from random sequences of the Pile dataset [22]. The Pile was used to train the Pythia models and includes OpenWebText, the training corpus for GPT-2. Testing models on data similar to their training data ensures a fair comparison and minimizes the impact of domain shifts on observed inference patterns.
| Parameters | Layers |
| Pythia (410M) | 24 |
| Pythia (1.4B) | 24 |
| Pythia (2.8B) | 32 |
| Pythia (6.9B) | 32 |
| Parameters | Layers |
| Small (124M) | 12 |
| Medium (355M) | 24 |
| Large (774M) | 36 |
@@
| Parameters | Layers |
| Phi 1 (1.3B) | 24 |
| Phi 1.5 (1.3B) | 24 |
| Phi 2 (2.7B) | 32 |
@@
To study the robustness and role of different model components at different depths, we employ a swapping intervention where we switch the execution order of a pair of adjacent layers in the model. Specifically, for a swap intervention at layer , we execute the transformer block (including the attention layer, MLP, and normalization) before executing block . We record the Kullback-Leibler (KL) divergence between the intervened and original models, measuring the difference in their output distributions, along with model-wise metrics such as loss, top-1 prediction accuracy, and prediction entropy. This intervention allows us to examine how the order of computation affects the model’s behavior and performance at different depths.
@@
To generate baselines for each layer swap experiment, we perform zero ablations on the corresponding layer while collecting the same metrics. The ablation preserves the swap ordering: for a swap ordering of 1-2-4-3-5, the ablation maintains 1-2-4-5. We opt for zero ablation as opposed to mean ablation, as proposed by [3], to maintain consistency with the swap order. Additionally, we perform attention-only and MLP-only ablations to study the specific roles of these components in the model’s inference process. By comparing the effects of layer swapping and ablation, we can gain insights into the importance and function of each component at different depths in the model.
@@ 4.1 Intervention Results
To study the robustness of language models, we apply our aforementioned drop and swap interventions to every layer of four GPT-2 models [55] and four Pythia [6]. In Figure 3, we report (1) the KL divergence between the prediction of the intervened model and the nominal model, (2) the fraction of predictions that are the same between the intervened model and the baseline model (denoted as relative accuracy), and (3) the change in entropy of the prediction between the intervened and baseline model for all interventions.
Our results show that intervening on the first layer is catastrophic for model performance. Specifically, dropping or swapping the first layer causes the model to have very high entropy predictions as opposed to causing a mode collapse on a constant token. In Pythia models, swapping the last layer with the second to last layer also has a similar catastrophic high-entropy effect, while GPT-2 models largely preserve their predictions. We further discuss our hypothesized role of the first and last layer in Section 5.1 and Section 5.4 respectively.
In contrast to the first and last layer interventions, the middle layers are remarkably robust to both deletion and minor order changes. When zooming in on the differences between the effect of swaps and drops for intermediate layers, we find that swapping adjacent layers is less harmful than ablating layers. These results match similar experiments performed on vision transformers [5]. We take this as evidence that certain operations within the forward pass are commutative, though further experimentation is required.
We suspect that GPT-2 exhibits greater robustness than Pythia because (1) GPT-2 models are trained with dropout, which likely increases redundancy and (2) GPT-2 models have fewer parameters per layer so a GPT-2 layer ablation removes fewer parameters than a Pythia ablation.
4.2 Why are Language Models Robust to Layer-Wise Interventions?
We suspect that the robustness of language models can be partially attributed to the presence of residual connections [65] in the transformer architecture, which lead to increased redundancy [28, 45]. While skip connections were initially introduced to mitigate the vanishing gradient problem, investigations of deep residual networks (ResNets) found that residual connections facilitate "ensembling" within networks [34, 65]. We hypothesize that, as is the case for ResNets, residual connections in language models allow gradient descent to form ensembles of relatively shallow computational sub-networks. In doing so, networks avoid strong dependencies on individual paths, thereby increasing their resilience to layer-wise interventions. This hypothesis is supported by recent observations of self-repair mechanisms [44, 58] that demonstrate the existence of parallel computational paths.
@@ 5 Stages of Inference Hypothesis
We now discuss and provide tentative evidence for our hypothesis of four universal stages of inference in transformer language models. Two caveats apply to all of the following subsections: First, the boundaries between stages are fuzzy, and in practice, more than one stage can occur simultaneously. Second, these stages represent aggregate computational patterns, but the processing of any specific kind of token is likely to undergo more individualized dynamics (e.g., factual recall [46, 49]).
@@ 5.1 Stage 1: Detokenization
Given the extreme sensitivity of the model to first-layer ablations, we infer that the first layer is not so much a normal layer as it is an extension of the embedding. This is especially true for the Pythia model family due to the use of parallel attention, which implies that the first MLP layer is only a function of the current token. Consequently, by ablating the first layer, the rest of the network is blind to the immediate context and is thrown off distribution.
Immediately after computing this extended embedding, evidence from the literature suggests that the model concatenates nearby tokens that are part of the same underlying word [9, 20] or entity [49] (e.g., a first and last name). This operation integrates local context to transform raw token representations into coherent entities. In this way, the input is “detokenized” [15, 31].
Previous work has shown the existence of neurons that activate for specific -grams [31, 66]. Of course, to accomplish this, there must be attention heads that copy nearby previous tokens into the current token’s residual stream. More generally, if early layers are integrating local context, we would predict that early-layer attention heads pay disproportionately more attention to nearby tokens than later attention heads.
To test this, we compute the fraction of attention paid to tokens within the previous five positions of the present token. As can be seen in Figure 5, attention is indeed more local in the early layers. Moreover, the output norm of the attention ratio is higher than the output norm of the MLP layers in the first few layers [54].
@@ 5.2 Stage 2: Feature Engineering
Building upon the locally contextualized representations from stage 1, we hypothesize that a second region of the model performs "feature engineering" to construct features that could be useful for making downstream predictions, either for the next token or for future tokens in the context. While the kinds of features will vary by token type, there are many results in the literature that localize intermediate feature construction to the early to middle layers via patching [70] and probing [2] experiments.
For example, the model editing literature suggests that the MLPs in this region are important for factual recall [46, 25, 49]. This region is where probing accuracy for spatial and temporal features drastically but smoothly improves [30]. The features produced in this stage influence downstream predictions, as evidenced by steering and patching experiments on sentiment [62], truth [41], and zero-shot function execution [63] that peak in effectiveness in stage two layers. The representations and features formed in this stage are increasingly abstract, transitioning from shallower syntactic features to richer semantic features [15, 68, 38].
@@
However, in this stage, the features are simply produced, rather than consumed to make a concrete prediction. To show this, we perform a logit lens experiment [50, 12] where we apply the unembedding to the residual stream after every layer to estimate a model’s intermediate prediction. We then compute the entropy of this intermediate prediction and the KL divergence with the final prediction. As can be seen in Figures 10 and 6, there is very little progress towards making an actual prediction. For that, we require the next two stages.

### v2

@@ Abstract
We investigate the robustness of Large Language Models (LLMs) to structural interventions by deleting and swapping adjacent layers during inference. Surprisingly, models retain 72–95% of their original top-1 prediction accuracy without any fine-tuning. We find that performance degradation is not uniform across layers: interventions to the early and final layers cause the most degradation, while the model is remarkably robust to dropping middle layers. This pattern of localized sensitivity motivates our hypothesis of four stages of inference, observed across diverse model families and sizes: (1) detokenization, where local context is integrated to lift raw token embeddings into higher-level representations; (2) feature engineering, where task- and entity-specific features are iteratively refined; (3) prediction ensembling, where hidden states are aggregated into plausible next-token predictions; and (4) residual sharpening, where irrelevant features are suppressed to finalize the output distribution. Synthesizing behavioral and mechanistic evidence, we provide a framework for interpreting depth-dependent computations in LLMs.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have exhibited remarkable reasoning capabilities, often attributed to increased scale [1]. Understanding these capabilities and mitigating associated risks [2, 3, 4] has motivated extensive research into their underlying mechanisms.
A bottom-up approach to interpretability, known as mechanistic interpretability, has explored the iterative inference hypothesis [5, 6], which posits that each transformer layer incrementally updates a token’s hidden state toward minimizing loss by progressively shaping the next-token distribution [7]. This is supported by self-repair [6], where later layers correct or mitigate errors introduced by earlier layers, and redundancy [8, 9], where multiple layers perform similar or overlapping computations to refine predictions.
However, this iterative view contrasts with the “circuit” hypothesis, which argues for clearly delineated, specialized roles for certain model components. This is supported by induction heads [10], successor heads [11], copy suppression mechanisms [12], and knowledge neurons [13], among other “universal” neurons [14, 15]. Whereas iterative inference suggests gradual refinement through overlapping computations, the strong circuit hypothesis implies distinct, modular computational pathways. Resolving this tension, specifically how specialized circuits integrate with iterative refinement processes, remains unclear. [10, 16].
Naturally, layer-wise phenomena in LLMs are also documented outside formal interpretability research and provide more evidence to existing interpretability findings. For example, while knowledge storage within mid-layer MLP neurons has been demonstrated [17], other non-interpretability work has found that fine-tuning predominantly affected the weights in the middle layers [18]. Quantization studies identified improved benchmark performance by retaining only low-rank MLP components from the middle to later layers [19]. Other works have noted a transition in activation sparsity from sparse to dense around mid-model depth [20, 15]. These behavioral findings, when integrated with mechanistic insights, suggest a layered computation structure not yet fully characterized.
To explore this structure, we perform layer-wise interventions—deleting individual layers or swapping adjacent ones (Figure 7)—to characterize their localized effects. Building on these insights, we analyze depth-wise roles and synthesize our findings with prior interpretability work to propose a four-phase framework that attempts to bridge the top-down and bottom-up view of computation in decoder-only LLMs.
Concretely, we hypothesize four depth-dependent stages: (1) detokenization, (2) feature engineering, (3) prediction ensembling, and (4) residual sharpening. In brief, early layers integrate local context to form coherent entities; middle layers iteratively construct features; later layers convert these features into next-token predictions via an ensemble of neurons; and final layers refine the output by suppressing noisy components. Figure 1 and Table 1 summarize these stages and their associated empirical signatures. We synthesize these findings with prior interpretability work [16] to suggest a depth-aligned computational structure in LLMs.
@@
1 | Detokenization | Integrate local context to transform raw token representations into coherent entities | Catastrophic sensitivity to deletion and swapping and attention-heavy computation. |
2 |
Feature
Engineering |
@@
Iteratively build feature representation depending on token context | Little progress made towards next token prediction, but significant increase in probing accuracy and patching importance. |
3 |
Prediction
Ensembling |
@@
Convert previously constructed semantic features into plausible next token predictions using an ensemble of model components | Prediction neurons appear and output distribution begins to narrow. |
4 |
Residual
Sharpening |
@@
Sharpen the next token distribution by eliminating obsolete features that add noise to the prediction |
Suppression neurons appear and output distribution narrows with a growing MLP-output norm
@@ 2 Experimental Protocol
To investigate the stages of inference in language models, we examine the Pythia [22], GPT-2 [23], Qwen 2.5 [24], LLaMA 3.2 [25], and Microsoft Phi [26, 27] model families, which range from 124M to 6.9B parameters (see Table 2). All families use decoder-only transformers but differ in their execution of attention and MLP components. Specifically, Pythia models execute attention and MLP layers in parallel. In contrast, GPT-2, Phi, and LLaMA models apply attention followed by an MLP sequentially. We preprocess weights identically across all models, folding in the layer norm, centering the unembedding weights, and centering the writing weights as described in Appendix A.11. Despite these architectural differences, most phenomena remain consistent across models, though we discuss drawbacks in Limitations 6.
@@
We evaluate all five model families on a corpus of one million tokens from random sequences of the Pile dataset [28], unless otherwise noted in the experiment.
Model Series | Size | Layers |
Pythia | 410M | 24 |
1.4B | 24 | |
2.8B | 32 | |
6.9B | 32 | |
GPT-2 | Small (124M) | 12 |
Medium (355M) | 24 | |
Large (774M) | 36 | |
@@
Model Series | Size | Layers |
Microsoft Phi | Phi-1 (1.3B) | 24 |
Phi-1.5 (1.3B) | 24 | |
Phi-2 (2.7B) | 32 | |
Llama 3.2 | 1B | 16 |
3B | 28 | |
Qwen 2.5 | 0.5B | 24 |
1.5B | 28 | |
3B | 36 |
@@
To study the robustness and role of different model components at different depths, we employ a swapping intervention where we switch the execution order of a pair of adjacent layers in the model. Specifically, for a swap intervention at layer , we execute the transformer block (including the attention layer, MLP, and normalization) before executing block . We record the Kullback-Leibler (KL) divergence between the intervened and original models output distribution, along with the loss, top-1 prediction accuracy, prediction entropy, and benchmark task performance. This intervention allows us to examine how the order of computation affects the model’s behavior and performance at different depths.
@@
To generate baselines for each layer swap experiment, we perform zero ablations on the corresponding layer while collecting the same metrics. The ablation preserves the swap ordering: for a swap ordering of 1-2-4-3-5, the ablation maintains 1-2-4-5. We opt for zero ablation as opposed to mean ablation, as proposed by [5], to maintain consistency with the swap order.
@@ 3 Robustness
(a) Layer Interventions Experiment
(b) CKA Similarity GPT2-XL
(c) CKA Similarity Pythia 2.8B
@@ 3.1 Intervention Results
We apply our aforementioned drop and swap interventions to every layer of four GPT-2 models [29] and four Pythia [22]. In Figure 2, we report (1) the KL divergence between the prediction of the intervened model and the nominal model, (2) the fraction of predictions that are the same between the intervened model and the baseline model (denoted as relative accuracy). We also report the performance on common benchmark tasks (HellaSwag[30], ARC-Easy[31] and LAMBADA[32]) for all models in Figure 9-10, which show a similar trend.
In contrast to the first and last layer interventions, the middle layers are remarkably robust to both deletion and minor order changes. When zooming in on the differences between the effect of swaps and drops for intermediate layers, we find that swapping adjacent layers is less harmful than ablating layers, matching a result in vision transformers [33]. We take this as an indication that certain operations within the forward pass are commutative, though further experimentation is required.
Intervening on the first layer is catastrophic for model performance for every model, regardless of size or model family. Specifically, dropping or swapping the first layer causes the model to have very high entropy predictions as opposed to causing a mode collapse on a constant token. In some models, swapping the last layer with the second-to-last layer also has a similar catastrophic high-entropy effect, while GPT-2 models largely preserve their predictions. This phenomenon motivates our study into the first few layers of the model, specifically the role paid by attention heads in these layers.
@@ 4 Stages of Inference Hypothesis
Motivated by the distinct phenomena at the first few and final few layers, we measured representational similarity across each layer output using Centered Kernel Analysis (CKA)[34, 35, 36]. This revealed a block-like structure across multiple models as shown in Figure 2. The existence of blocks reflects the robustness observed in the layer-wise intervention. Furthermore, the depth-dependent phase structure indicates that a shared computation motif across adjacent layers occurs in stages.
@@ 4.1 Stage 1: Detokenization
Given the extreme sensitivity of the model to first-layer ablations, we infer that the first layer is not a normal layer, but rather an extension of the embedding. Uniquely, the first layer is the layer that moves from the embedding basis to that of the transformer’s residual stream. It is only a function of the current token. Consequently, by ablating the first layer, the rest of the network is blind to the instant context and is thrown off distribution. Immediately after computing this extended embedding, evidence from the literature suggests that the model concatenates nearby tokens that are part of the same underlying word [37, 38] or entity [39] (e.g., a first and last name). This operation integrates local context to transform raw token representations into coherent entities. In this way, the input is “detokenized” [40, 41]. Previous work has shown the existence of neurons that activate for specific -grams [41, 15]. Of course, to accomplish this, there must be attention heads that copy nearby previous tokens into the current token’s residual stream.
Subjoiner Heads
To further examine this detokenization mechanism, we investigated attention heads responsible for constructing multi-token words, known as subjoiner heads [38]. These heads help capture the context of a token for appropriate prediction, thus contributing to the detokenization process. We constructed a dataset with two classes: each consisting of 16 tokens, where in one class, the final 4 tokens form a word. Our analysis identified specific heads in the early layers of models that contribute solely to the construction of these multi-token words. As illustrated in Figure 3(b), layer 2 head 5 of Pythia 2.8B moves information from earlier tokens to the final token of the word. The attention heads exhibit a consistent pattern, where attention decreases as tokens approach the final word. Specifically, the final token of the word attends most strongly to the first token, a feature absent in the baseline. This suggests at least one of many mechanisms by which models integrate local context, occurring at higher density in the first half of the models.
Local Attention
If early layers indeed specialize in integrating local context, then we would expect attention heads in these layers to disproportionately focus on tokens close to the current position. To investigate this hypothesis, we measure the fraction of attention that each token directs toward preceding tokens at varying distances. As shown in Figure 3, attention heads in early layers are strongly biased towards nearby tokens, with attention becoming progressively less localized in deeper layers.
@@ 4.2 Stage 2: Feature Engineering
After integrating local context in the early layers—e.g., stitching together subword tokens and forming short-range dependencies—the model must begin converting those localized representations into more semantically meaningful features. We hypothesize that this marks the beginning of a feature engineering stage, in which the model constructs intermediate features that encode abstract properties useful for downstream prediction.
Prior work provides indirect support for this idea. Model editing studies suggest that factual information is stored in mid-layer MLPs [17, 42, 39], while probing experiments have found that intermediate layers encode features related to sentiment [43], truth [44], and temporal structure [45]. These studies typically show that probing accuracy rises through the early layers, peaks near the midpoint, and then declines, suggesting that features are constructed and later transformed or compressed. Related work also observes a shift from syntactic to semantic representations with depth [40, 46].
WiC Probing
To illustrate this pattern, we train linear probes to detect context-dependent word meaning using the WiC (Word-in-Context) task [47, 48]. For instance, given two sentences containing the word bank, the task is to classify whether it is used with the same meaning. Examples include distinguishing “the river bank” from “the robbed bank,” where the same word has different meanings depending on the context. We apply this probe at each layer of the model, using the hidden state of the target word in context. As shown in Figure 4 (left), the accuracy of the probe increases through the early layers, peaks in the middle of the model, and then decreases, supporting the hypothesis that semantic features are most linearly accessible in the intermediate layers. We extend the observation across model families and sizes in Figure 14.
@@
While these results suggest that intermediate representations encode semantic information, it remains unclear whether such features contribute to prediction at this stage. To investigate this, we apply the logit lens [49, 50], which projects the residual stream at each layer into the output vocabulary space using the model’s unembedding matrix. This provides a layer-wise estimate of the model’s next-token distribution.
We compute both the entropy of the intermediate predictions and their KL divergence from the model output. As shown in Figure 4 (right), entropy remains high and KL divergence low throughout the early and middle layers. In other words, while meaningful features appear to be present in the residual stream at this stage, the model’s output distribution remains high in entropy, indicating that these features have not yet been consolidated into confident next-token predictions. Bridging this gap requires a mechanism that selectively retains information from relevant features while filtering out irrelevant ones, thereby reducing uncertainty in the output distribution.